# A short character sum in $\mathbb{F}_{p^{3}}$

### v1

@@ 6. Application to Character Sums of Binary cubic form
$<$ $exp\left(c_{k}\frac{\log p}{\log\log p}\right).\frac{|T|^{2}}{p^{2}}\sum_{\chi_{1},\chi_{2},\chi_{3}}\left|\chi_{1}(t)\chi_{2}(t)\chi_{3}(t)\right|^{2}\left|\sum_{x\in I,y\in J}\chi_{1}(x-\lambda_{1}y)\chi_{2}(x-\omega_{2}y)\chi_{3}(x-\omega_{3}y)\right|^{2}$
$=$ $exp\left(c.\frac{\log p}{\log\log p}\right).|T|^{2}E(R,\Delta)$
@@
(27) $Q<=exp\left(c.\frac{\log p}{\log\log p}\right).|T|^{2}E(R,\Delta)$
@@
(28) $\begin{cases}t(x-\lambda_{1})=t^{\prime}(x^{\prime}-\lambda_{1}y^{\prime})\bmod{F_{p}}\\ t(x-\omega_{2})=t^{\prime}(x^{\prime}-\omega_{2}y^{\prime})\bmod{F_{p}}\\ t(x-\omega_{3})=t^{\prime}(x^{\prime}-\omega_{3}y^{\prime})\bmod{F_{p}}\\ \end{cases}$
@@
(29) $\begin{cases}tx=t^{\prime}x^{\prime}\bmod{p}\\ ty=t^{\prime}y^{\prime}\bmod{p}\\ \end{cases}$
which implies $xy^{\prime}=x^{\prime}y\bmod{p}$ and number of solution set of it is bounded by $E(I,J)\lesssim(\log p).|I||J|$ where $|I|,|J|<p$. Once $x.x^{\prime},y,y^{\prime}$ are fixed, the number of solutions $(t,t^{\prime})$ in(28) is at most $(p-1)$. Hence, the number of solutions of equation(28) is bounded by $p(\log p)|I||J|$. So,
@@
Using this in 27 we get
(30) $Q\leq exp\left(c\frac{\log p}{\log\log p}\right)p(\log p)|T|^{2}.|R|.$
@@
$E(R,T,S)$ $\leq exp\left(c\frac{\log p}{\log\log p}\right)p^{5/2k}R^{2-1/k}$
$\leq exp\left(c\frac{\log p}{\log\log p}\right)p^{3/k}(|I||J|)^{2-1/k}$
$\leq exp\left(c\frac{\log p}{\log\log p}\right)|K|^{2}(|I||J|)^{2-2\kappa/3}$
@@
Case 3:
Suppose $f(x,y)$ is completely reducible in distinct factors:
@@
We shall proceed by the same way by replacing $\mathbb{F}_{p^{2}}$ by $\mathbb{F}_{p}^{3}$. Denote: $I_{0}=[1,p^{3/8}]$ and $K$ in the previous way and translate $(x,y)$ by $(stx_{0},sty_{0})$ we obtain a character sum estimate
(32) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}x\in I,y\in J\\ x_{0},y_{0}\in I_{0}\\ s\in K\end{subarray}}\left|\sum_{t\in K}\chi\left(\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{2}y_{0})}\right)\left(t+\frac{x-\lambda_{3}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\right)\right|$
@@
(33) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2},z_{3})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{3})\right|$
In accordance with previous case, if we apply Holder’s inequality twice here we analogously get
(34) $\frac{1}{|K|^{2}|I_{0}|^{2}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2},z_{3})\right)^{1-\frac{1}{k}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2},z_{3})^{2}\right)^{\frac{1}{2k}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{3})\right|^{2k}\right)^{\frac{1}{2k}}$
The first factor of the above product is trivially bounded by $(|I||J||I_{0}||I_{0}||K|)^{1-\frac{1}{k}}$. By suitably applying Weil bound[12] the third factor is estimated as Lemma 13
$\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{3}))\right|^{2k}\right)^{\frac{1}{2k}}\leq k|K|^{1/2}p^{3/2k}+2|K|p^{3/4k}.$
Here,
$\sum w(z_{1},z_{2},z_{3})^{2}=E(R,T,S)$
where,
$R=\{(x-\lambda_{1}y,x-\lambda_{2}y,x-\lambda_{3}y):x\in I,y\in J\}$
$T=\{(x_{0}-\lambda_{1}y_{0},x_{0}-\lambda_{2}y_{0},x_{0}-\lambda_{3}y_{0}):x_{0},y_{0}\in I_{0}\}$
@@
So, it is suffices to bound the multiplicative energy of the set $(R,T,S)$ by
(35) $\sum_{z_{1},z_{2},z_{3}\in\mathbb{F}_{p}}w(z_{1},z_{2},z_{3})^{2}<|I|^{2}|J|^{2}|K|^{2}p^{-\tau}$
for some $\tau>0$. And it can be calculated analogously to the previous case.
Case 4:
If $f(x,y)$ is completely factorizable in $\mathbb{F}_{p}$, but not all roots are distinct from each other. In this case, the polynomial $f(x,y)$ should look like
$f(x,y)=(x-\lambda_{1}y)(x-\lambda_{2}y)^{2}.$
By proceeding as the above cases and translating $(x,y)$ suitably we obtain:
(36) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})^{2})\right|$
where,
(37) $w(z_{1},z_{2})=\left|\left\{(x,y,x_{0},y_{0},s)\in I\times J\times I_{0}\times I_{0}\times K:z_{1}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})},z_{2}=\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})}\right\}\right|.\$
Applying Holder’s twice as usual we obtain
(38) $\frac{1}{|K|^{2}|I_{0}|^{2}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2})\right)^{1-\frac{1}{k}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2})^{2}\right)^{\frac{1}{2k}}\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})^{2})\right|^{2k}\right)^{\frac{1}{2k}}.$
Now, the first factor is bounded trivially and the last factor is bounded by $k|K|^{1/2}p^{2/2k}+2|K|p^{2/4k}=k|K|^{1/2}p^{1/1k}+2|K|p^{1/2k}$ due to Weil[12]. The proof is almost similar to Lemma13. The required upper bound of the middle one is
$\sum_{z_{1},z_{2}}w(z_{1},z_{2})^{2}<|I|^{2}|J|^{2}|K|^{2}p^{-\tau}.$
for some $\tau>0$. Here, if we assume $|I|,|J|<p$. Let
$R=\{(x-\lambda_{1}y,x-\lambda_{2}y):x\in I,y\in J\},$
$T=\{(x_{0}-\lambda_{1}y_{0},x_{0}-\lambda_{2}y_{0}):x_{0},y_{0}\in I_{0}\},$
$S=\{(s,s):s\in K\}.$
Considered as a subset of $\mathbb{F}_{p}^{*}\times\mathbb{F}_{p}^{*}.$ So, the required bound is equivalent with
$E(R,T,S)\leq p^{-\tau}|I|^{2}|J|^{2}|K|^{2}$
which can be established in the similar manner as Case 2. Again, fixing $|K|=p^{3/2k}$ and combining the above three estimates in (38), we obtain a nontrivial upper bound $|I||J|p^{-\tau/2k}$ for (36).
7. Appendix
First, we state Weil’s bound for character sums.
Theorem 12.
@@ 7. Appendix
See [Chapter $11$,[13]]. ∎
We shall use the above estimate to find a bound for 21
@@
Let $\chi$ be a nontrivial multiplicative character of $\mathbb{F}^{*}$ of order $d>1$. Then
$\left(\sum\limits_{\begin{subarray}{c}a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\\ c\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+a)(t^{2}+bt+c))\right|^{2k}\right)^{\frac{1}{2k}}\leq r|K|^{\frac{1}{2}}p^{\frac{3}{2r}}+|K|p^{\frac{3}{4r}}$
@@
$\sum\limits_{\begin{subarray}{c}a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\\ c\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+a)(t^{2}+bt+c))\right|^{2k}$
$=$ $\sum_{t_{1},t_{2},\ldots,t_{2r}\in K}\sum\limits_{\begin{subarray}{c}a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\\ c\in\mathbb{F}_{p}\end{subarray}}\prod_{i=1}^{2r}\chi_{i}((t_{i}+a)(t_{i}^{2}+bt_{i}+c))$
[…]
$\leq$ $\sum_{t_{1},t_{2},\ldots,t_{2r}\in K}\left(\sum_{a\in\mathbb{F}_{p}}\chi\left(\prod_{i=1}^{r}(t_{i}+a)\prod_{i=r+1}^{2r}(t_{i}+a)^{p-2}\right)\right)\left(\sum_{\omega\in\mathbb{F}_{p^{2}}}\chi\left(N\left(\prod_{i=1}^{r}(t_{i}+\omega)\prod_{i=r+1}^{2r}(t_{i}+\omega)^{p-2}\right)\right)\right)$
Now, we shall collect the 2r-tuples $(t_{1},t_{2},\ldots,t_{2r})\in K^{2r}$ where it violates Weil’s condition, which means that the polynomial function inside the character sum is a $d-$th power. We shall estimate character sum in the above region by trivial estimate, and in the rest of the region we shall apply the Weil bound [12].
Suppose in the above character sums $\{t_{i}\}_{1}^{r}$ are different from $\{t_{i}\}_{r+1}^{2r}$ and the polynomial inside is a $d-$th power, then it will violet the Weil’s condition. One can observe that the number of tuples $(t_{1},\ldots,t_{2r})\in K^{2r}$ where this happens is $|K|^{\frac{r}{d}}|K|^{\frac{r}{d}}$. Next, if one of the elements of the set $\{t_{i}\}_{1}^{r}$ is the same as one element of $\{t_{i}\}_{r+1}^{2r}$ then we need to choose the one element from both sets and count the number of other $t_{i}$s that form the polynomial function inside to be a $d-$th power. Here we obtain that there are at most $(^{r}C_{1})^{2}|K|^{\frac{r-1}{d}}|K|^{\frac{r-1}{d}}|K|$ points of the form $(t_{1},t_{2},\ldots,t_{2r})\in K^{2r}$ where Weil’s condition violets. Similarly, if two of the elements are same in both sets $\{t_{i}\}_{1}^{r}$ and $\{t_{i}\}_{r+1}^{2r}$ then the number of points$(t_{1},\dots,t_{2r})\in K^{2r}$ where the polynomials inside will be a $d-$th power will be bounded by $(^{r}C_{2})^{2}|K|^{\frac{r-2}{d}}|K|^{\frac{r-2}{d}}|K|^{2}$. And so on.
Hence, the total number of points in $K^{2r}$, where the polynomials are $d-$th powers, are bounded by
$\sum_{i=0}^{r}(^{r}C_{i})^{2}|K|^{\frac{2i}{d}}|K|^{r-i}$
[…]
$\leq$ $\sum_{i=0}^{r}(^{r}C_{i})^{2}|K|^{r}$
$\leq$ $(^{2r}C_{r})|K|^{r}\leq r^{2r}|K|^{r}$
We shall take the trivial bound for these bad points (where we can’t apply Theorem 12) and we get $r^{2r}|K|^{r}p^{3}$. Other than the above, all points are good points( i.e. they satisfy Weil’s condition). So in this case, we can apply Theorem [12] to both the sums, and we get an upper bound of size $6r|K|^{2r}p^{\frac{3}{2}}$. Finally we get
$\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t^{2}+at+b))\right|^{2r}\leq r^{2r}|K|^{r}p^{3}+6r|K|^{2r}p^{\frac{3}{2}}$
So,
$\left(\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ a\in\mathbb{F}_{p}\\ b\in\mathbb{F}_{p}\end{subarray}}\left|\sum_{t\in K}\chi((t+z_{1})(t^{2}+at+b))\right|^{2k}\right)^{\frac{1}{2k}}\leq r|K|^{\frac{1}{2}}p^{\frac{3}{2r}}+6|K|p^{\frac{3}{4r}}.$
This was the required estimate. ∎
Remark. Other "Weil type estimates" of this paper can be proved similar to the above.

### v3

@@ 5. Application to character sums of binary cubic form
$<$ $\exp\left(c_{k}\frac{\log p}{\log\log p}\right).\frac{|T|^{2}}{p^{2}}\sum_{\chi_{1},\chi_{2},\chi_{3}}\left|\chi_{1}(t)\chi_{2}(t)\chi_{3}(t)\right|^{2}\left|\sum_{x\in I,y\in J}\chi_{1}(x-\lambda_{1}y)\chi_{2}(x-\omega_{2}y)\chi_{3}(x-\omega_{3}y)\right|^{2}$
$=$ $\exp\left(c.\frac{\log p}{\log\log p}\right).|T|^{2}E(R,\Delta)$
@@
(26) $Q\leq\exp\left(c.\frac{\log p}{\log\log p}\right).|T|^{2}E(R,\Delta)$
@@
[…]
@@
[…]
which implies $xy^{\prime}=x^{\prime}y\bmod{p}$ and number of solution set of it is bounded by $E(I,J)\lesssim(\log p).|I||J|$ where $|I|,|J|<p$. Once $x.x^{\prime},y,y^{\prime}$ are fixed, the number of solutions $(t,t^{\prime})$ in(27) is at most (p1)(p-1). Hence, the number of solutions of equation(27) is bounded by $p(\log p)|I||J|$. So,
@@
Using this in 26 we get
(29) $Q\leq\exp\left(c\frac{\log p}{\log\log p}\right)p(\log p)|T|^{2}.|R|.$
@@
$E(R,T,S)$ $\leq\exp\left(c\frac{\log p}{\log\log p}\right)p^{5/2k}R^{2-1/k}$
$\leq\exp\left(c\frac{\log p}{\log\log p}\right)p^{3/k}(|I||J|)^{2-1/k}$
$\leq\exp\left(c\frac{\log p}{\log\log p}\right)|K|^{2}(|I||J|)^{2-2\kappa/3}$
@@
Case 3. $f(x,y)$ splits completely in $\mathbb{F}_{p}$ with distinct linear factors. Suppose
@@
The same argument applies over $\mathbb{F}_{p}\times\mathbb{F}_{p}\times\mathbb{F}_{p}$ in place of $\mathbb{F}_{p}\times\mathbb{F}_{p^{2}}$. Here $I_{0}$ and $K$ are the same as before. Translating $(x,y)$ by $(stx_{0},sty_{0})$ we obtain a character sum estimate
(31) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}x\in I,y\in J\\ x_{0},y_{0}\in I_{0}\\ s\in K\end{subarray}}\left|\sum_{t\in K}\chi\left(\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\left(t+\frac{x-\lambda_{1}y}{s(x_{0}-\lambda_{2}y_{0})}\right)\left(t+\frac{x-\lambda_{3}y}{s(x_{0}-\lambda_{1}y_{0})}\right)\right)\right|.$
@@
(32) $\frac{1}{|K|^{2}|I_{0}|^{2}}\sum\limits_{\begin{subarray}{c}z_{1}\in\mathbb{F}_{p}\\ z_{2}\in\mathbb{F}_{p}\\ z_{3}\in\mathbb{F}_{p}\end{subarray}}w(z_{1},z_{2},z_{3})\left|\sum_{t\in K}\chi((t+z_{1})(t+z_{2})(t+z_{3})\right|.$
rf
After applying Hölder’s inequality and the Weil bound (a variant of Lemma 11), it remains to show that
$E(R,T,S)\leq p^{-\tau}|I|^{2}|J|^{2}|K|^{2},$
where
$R$ $=\{(x-\lambda_{1}y,\,x-\lambda_{2}y,\,x-\lambda_{3}y):x\in I,\,y\in J\},$
$T$ $=\{(x_{0}-\lambda_{1}y_{0},\,x_{0}-\lambda_{2}y_{0},\,x_{0}-\lambda_{3}y_{0}):x_{0},y_{0}\in I_{0}\},$
@@
The estimation of $E(R,T,S)$ proceeds exactly as in Case 2, yielding the same bound as required. ∎
Appendix
We recall the classical Weil bound for multiplicative character sums, which plays a key role in our estimates.
Theorem 9 (Weil).
@@
See [16, Chapter 11]. ∎
A more general theorem in this regard is the following
Theorem 10 (Weil bound for multiplicative characters on $\mathbb{F}_{p^{d}}$).
Let $q=p^{n}$ be a power of a prime and let $\chi:\mathbb{F}_{q}^{\times}\to\mathbb{C}^{\times}$ be a nontrivial multiplicative character of order $d$. Extend $\chi$ to $\mathbb{F}_{q}$ by defining $\chi(0)=0$. Let $f(x)\in\mathbb{F}_{q}[x]$ be a polynomial which has $m\geq 1$ distinct root such that and is not an $d$-th power in $\overline{\mathbb{F}}_{q}[x]$. Then
$\left|\sum_{x\in\mathbb{F}_{q}}\chi\big(f(x)\big)\right|\leq(m-1)q^{1/2}.$
Proof.
See the proof of [20, Theorem 2C’, p.43].
The following lemma provides the key averaged estimate used in the main argument.
@@
Let $\chi$ be a nontrivial multiplicative character of $\mathbb{F}_{p}^{*}$ of order $d>1$. Then for any integer $r\geq 1$,
$\left(\sum_{a,b,c\in\mathbb{F}_{p}}\Big|\sum_{t\in K}\chi\big((t+a)(t^{2}+bt+c)\big)\Big|^{2r}\right)^{\!\!\frac{1}{2r}}\ll r|K|^{1/2}p^{\frac{3}{2r}}+|K|p^{\frac{3}{4r}}.$
@@
Expanding the $2r$-th power, we obtain
$\sum_{a,b,c}\!\Big|\sum_{t\in K}\chi((t+a)(t^{2}+bt+c))\Big|^{2r}=\!\sum_{t_{1},\dots,t_{2r}\in K}\!\sum_{a,b,c}\prod_{i=1}^{2r}\chi_{i}\!\big((t_{i}+a)(t_{i}^{2}+bt_{i}+c)\big),$
where $\chi_{i}=\chi$ for $1\leq i\leq r$ and $\chi_{i}=\overline{\chi}$ for $r<i\leq 2r$. The inner sum factors into two character sums—one over $\mathbb{F}_{p}$ and one over its quadratic extension $\mathbb{F}_{p^{2}}$ as follows
$\sum_{t_{1},t_{2},\ldots,t_{2r}\in K}\left|\sum_{a\in\mathbb{F}_{p}}\chi\left(\prod_{i=1}^{r}(t_{i}+a)\prod_{i=r+1}^{2r}(t_{i}+a)^{p-2}\right)\right|\times\left|\sum_{\omega\in\mathbb{F}_{p^{2}}}\chi\left(N\left(\prod_{i=1}^{r}(t_{i}+\omega)\prod_{i=r+1}^{2r}(t_{i}+\omega)^{p-2}\right)\right)\right|.$
Step 1. Counting exceptional tuples. We call a tuple $(t_{1},\dots,t_{2r})\in K^{2r}$ bad if the polynomial inside the character becomes a $d$-th power in $\mathbb{F}_{p}[X]$; in this case Weil’s bound does not apply. Now in any bad tuple every element occurs at least twice. Consequently, the number of such bad tuples is bounded by
$\{\text{bad tuples}\}\leq|K|^{r}.$
For these tuples, we use the trivial estimate for the character sums, giving a contribution $\leq|K|^{r}p^{3}$.
Step 2. Applying Weil’s bound. For all other (good) tuples, the conditions of Theorem 9 are satisfied, and each of the two character sums are bounded by $O(r\,p^{1/2})$ and O(rp)O(r\,p) respectively. Hence their combined contribution is $\ll r^{2}|K|^{2r}p^{3/2}$.
Step 3. Combining both cases. Summing these contributions we obtain
$\sum_{a,b,c}\Big|\sum_{t\in K}\chi((t+a)(t^{2}+bt+c))\Big|^{2r}\ll|K|^{r}p^{3}+r^{2}|K|^{2r}p^{3/2}.$
Taking the $2r$-th root yields the desired bound. All implied constants are absolute. ∎
Remark 3. Other Weil type estimates of this paper can be proved similarly.